Direct Method of Support Department Cost Allocation

Valron Company has two support departments, Human Resources and General Factory, and two producing departments, Fabricating and Assembly.


Support Departments
Producing Departments
Human
Resources
General
Factory
Fabricating Assembly
Direct costs $160,000 $340,000 $114,600 $93,000
Normal activity:
Number of employees 60 80 170
Square footage 1,000 5,700 13,300

The costs of the Human Resources Department are allocated on the basis of number of employees, and the costs of General Factory are allocated on the basis of square footage. Valron Company uses the direct method of support department cost allocation.

Required:

1. Calculate the allocation ratios for the four departments using the direct method. If an amount is zero, enter "0". Round your answer to the nearest cent.

Proportion of Driver Used by
Human Resources General Factory Fabricating Assembly
Human Resources
General Factory

2. Using the direct method, allocate the costs of the Human Resources and General Factory departments to the Fabricating and Assembly departments. If an amount is zero, enter "0".

Support Departments Producing Departments
Human Resources General Factory Fabricating Assembly
Direct costs $ $ $ $
Allocate:
  Human Resources            
  General Factory            
Total after allocation $ $ $ $
Answer #1

Answer

  • Requirement 1

Proportion of Driver Used by

Human Resources

General Factory

Fabricating

Assembly

Human Resources

                        -  

                                  -  

                         0.3200

                       0.6800

General Factory

                        -  

                                  -  

                                 0.3000

                                0.7000

--Working

Proportion of Driver Used by

Human Resources

General Factory

Fabricating

Assembly

Human Resources

0

0

=80/(80+170)

=170/(80+170)

General Factory

0

0

=5700/(5700+13300)

=13300/(5700+13300)

  • Requirement 2

Support Departments

Producing Departments

Human Resources

General Factory

Fabricating

Assembly

Direct costs

$160,000

$340,000

$114,600

$93,000

Allocate:

Human Resources

($160,000)

$51,200

$108,800

General Factory

$0

($340,000)

$102,000

$238,000

Total after allocation

$0

$0

$267,800

$439,800

--Working

Support Departments

Producing Departments

Human Resources

General Factory

Fabricating

Assembly

Direct costs

160000

340000

114600

93000

Allocate:

Human Resources

-160000

=160000*0.32

=160000*0.68

General Factory

-340000

=340000*0.3

=340000*0.7